OpenStudy (anonymous):

give the general formula for cos(theta)=-(square root 3)/2. list six solutions

OpenStudy (anonymous):

Find the general solution and you can churn out as many solutions as you want.

OpenStudy (anonymous):

i understand that theta = -\[\sqrt{3}/2\] at 5pi/6 and 7pi/6 but im stumped from there

OpenStudy (anonymous):

Add \(2\pi\)

OpenStudy (anonymous):

its asking for 1 formula, but the only way i know how to do it is \[\theta=(5\pi/6 + 2\pi k) \] and \[\theta=(7\pi/6 + 2\pi k ) \] where k is an integer. how do i make it 1 formula?

OpenStudy (anonymous):

Hmmmm, I'm not completely sure.

OpenStudy (anonymous):

Usually you can combine answers when they are \(\pi\) away.

OpenStudy (anonymous):

thats the prob im having, its a negative so cos can only b neg on Quadrant 2 & 3 so its cant b ...

OpenStudy (anonymous):

There isn't a nice way of combining them.

OpenStudy (anonymous):

However, if you go halfway between the angles..

OpenStudy (anonymous):

You could do this.\[ \frac{(6\pm 1)\pi }{6}+2\pi k \]

OpenStudy (anonymous):

that makes it way cleaner. much appreciated . Semper Fi.